Subject: Minutes of May 11 3.5m User's Committee meeting

From: strauss@astro.Princeton.EDU

Submitted: Wed, 13 May 1998 13:45:53 -0400 (EDT)

Message number: 277 (previous: 276, next: 278 up: Index)

	Apache Point Observatory 3.5m User's Committee Meeting
****************May 11, 1998**********************

Attending: Jon Holtzman, Michael Strauss, Bruce Gillespie, Ed
Turner, Jeff Brown, Alan Uomoto, Chris Stubbs, Ed Kibblewhite

  Status of Instruments
  New Dust Policy
  Recruitment Status at the Site
  Status of T1 line from Las Cruces

***********Status of Instruments**************

Gillespie: SPICAM is currently down; there was a shutter failure a
week ago.  There have been problems with the shutter for a while, so
equipment was bought to replace it with a more robust system, and this
is being done now during bright time.  The opportunity presents itself
to do some upgrades.  New microcode will be put in place which will
have enhanced abilities:
  The ability to abort in the middle of drift scan
  Larger range of drift-scanning speeds
  Faster clearing

  Should be all in place in a week, or less.

   Non-uniform focus has been noticed across the red chip, implying
that perhaps the chip has been mis-aligned.  We don't know if this is
a new or old problem. 
   There have been some recent problems with the interface between
Remark and Mirella code; with Lupton on site, some of these may be
straightened up.  Another problem: recent small "fixes" to the MC code
have caused FITS image bits to be corrupted; the problem has been
tracked down, and people were able to fix the one night of corrupted
data (these data were repairable).  We need to have tighter control of
software changes; the real problem is that no one on-site knew that
these changes had actually happened.  We discussed the possibility of
using CVS system for management of software changes, or simply
screening all such changes through Craig Loomis.

  Medium-resolution grating: a few people have reported that these
gratings were not put in place when they were requested.  To make sure
the proper gratings are put in, indicate this in your proposal, and
make sure to let the people on the mountain know of your grating needs
at least 24 hours in advance via e-mail to,
just like we do for filter configuration changes or confirmation.

   Holtzman: There are "talking" plans for several upgrades:
procurement and installation of a molecular hydrogen continuum filter
(which turns out to be quite expensive), an occulting spot, and the
ability to do multiple readouts.  Rauscher and Loewenstein indicate
that the multiple readouts should be fairly straightforward to
implement.  There is also need to do some mechanical work on the
instrument: maintenance of moving parts, cleaning of getter, etc.,
perhaps during July shutdown.  It is essential to get one of the
original builders of GRIM on the mountain for this work.

  Ed Kibblewhite: Visitor Instrument: Mid-IR (3-5 microns) 1Kx1K
camera built by Goddard; with fast read-out and real-time
shift-and-add (i.e., real-time speckle imaging), to get
diffraction-limited images over a few arcminutes.  They would like to
use ARC to demonstrate the technology for NGST; in exchange, the ARC
user's community would have access to this instrument for what might
be an extended period.  They have a test run at APO next month.

   Work is going ahead; the instrument is scheduled to come to APO
sometime in June, with installation during the shutdown in July.
Commissioning plan is being done by some combination of Roger Hildebrand
(, Lew Hobbs (,
and Doug Duncan (  Send them e-mail if you
are interested in taking part in commissioning or to provide input to the
testing requirements.

  Kibblewhite: I have funding from NSF, and will be back in the game
for ChAOS later this year.

  The guider is now fully commissioned.  Can it be used for real-time
focus of telescope?  The problem is that as it rotates around, it
moves in and out of the focal plane of the telescope.  This makes it
difficult to use it for focus.  There is a code to do an automated
focus ("gfocus") on the guider.  We just need to spend a large amount
of time to map out relative focus of guider and instrument as a
function of rotator angle, altitude of telescope, and telescope
temperature.  A good project for an interested user!

  Still at high priority; rotation of the tertiary, being worked by
UWashington under C. Stubbs direction.

***************Dust Policy************************

See apo35-general #275 for a new policy, which also begins the
coordination of site operations between 3.5m and the SDSS 2.5m and
Monitor Telescopes.  The original cutoff was designed to be a warning
to users that a dust storm was on its way.  But the appropriate
threshold for this can be a function of time of year.  Moreover, even
serious dust events (10,000's of counts) are not serious, if you clean
the mirror the following morning, *as long as it didn't get wet*!  (If
it does get wet, there is a real danger of the dust cementing to the
mirror).  Pollen is a particular problem this time of year, and it is
possible that it is not very well measured by the dust counter.
Pollen seems to be effectively removed by the standard mirror cleaning
procedure, although there is always the worry that some organic sticky
residue is left behind.
  Jim Crocker has suggested a spray-on stripping lacquer for cleaning
mirrors, which is used successfully at ESO.  New ways of washing
mirrors are also being looked at.

Current fan and filter system for drawing air into dome during the day
seems also to reduce dust around the telescope during the night; it
may affect the seeing and so should be used only when dust levels
start becoming threateningly high.

Stubbs: Throughput measurements off of polar fields are underway.


Temporary summer observing specialist from UW will be arriving soon to
replace Karen Gloria during her summer vacation.  Charles Corson's
replacement is also close to being on board.

*************T1 link****************************
It has been suggested to use the existing T1 line to Las Cruces to
connect to ISDN line; this could also be used for telephone and video
conferencing.  This could be used as a backup when regular internet
goes down.  The protocols need to be developed to make sure that the
telephone and video conferencing uses don't interfere with remote
observing.  ESO and Mauna Kea use video conferencing for remote
observing quite successfully; a possibility for us as well.

  Note: if the internet goes down, the observing specialist is not
expected to go into service observing mode!   There have been cases
where the remote observer has expected the observing specialist to
finish the scientific program of the night, after the internet
connection failed.

  The summer shutdown is scheduled for July 6-31 (see apo35-general

  Last month's minutes are approved.

  Next meeting, Monday, June 8, 12:30 PM

APO APO APO APO APO  Apache Point Observatory 3.5m  APO APO APO
APO  This is message 277 in the apo35-general archive. You can find
APO  the archive on
APO  To join/leave the list, send mail to
APO  To post a message, mail it to